# Clines and the Analytic Structure of Black Hole Perturbations

### v1

@@ 1 Introduction
Heun differential equations appear across physics—from quantum mechanics and condensed matter systems to perturbations in curved spacetimes—encoding field dynamics through their singularities and connection structure. In black hole physics, both in four and higher dimensions, Heun equations are particularly relevant to the study of quantum corrections to black hole entropy [Arnaudo:2024bbd, Arnaudo:2025btb], scattering processes and greybody factors [Cvetic:1997uw, Cvetic:2009jn, Jorge:2014kra], quasinormal modes [Berti:2009kk, Fioravanti:2021dce, Hatsuda:2021gtn, Bianchi:2021mft, Noda:2022zgk, Lei:2023mqx, Silva:2025khf, Chen:2025sbz], and tidal Love numbers [Kol:2011vg, Hui:2020xxx, Hui:2021vcv, Hui:2022vbh, Rodriguez:2023xjd, Perry:2023wmm, Charalambous:2023jgq, Perry:2024vwz, Glazer:2024eyi]. Beyond this context, similar techniques have found applications in holography, particularly in the study of renormalization group flows [Caceres:2021fuw, Caceres:2022smh], correlation functions [Gutperle:2025bzr] and deformations of the boundary conformal field theory [Karch:2025hof].
Over the last few years, the Heun connection formula program has achieved extraordinary success in describing black hole perturbations, including scattering processes [Bonelli:2021uvf, Bonelli:2022ten]. The method addresses the connection problem — which involves establishing the matrices that connect Frobenius series solutions around distinct singular points — for the local solutions of the four-point Heun equation, employing the so-called AGT correspondence introduced by Alday, Gaiotto, and Tachikawa [Alday:2009aq]. A central aspect is the use of conformal blocks provided by this duality to construct explicit connection formulas between the local solutions of the Heun equation [Bonelli:2022ten]. This approach has also been generalized to the confluent limit of the Heun equation [Bonelli:2021uvf, Bonelli:2022ten, Consoli:2022eey, Aminov:2023jve, Jia:2024zes, Gutperle:2025bzr]. An alternative approach for deriving connection formulas is based on expressing them as ratios of Wronskians constructed from independent solutions of the differential equation. This method, grounded in the analytic continuation properties of linear differential equations, allows one to extract connection coefficients without explicitly matching local Frobenius series. As demonstrated by Hatsuda, Kimura, Noda, and Motohashi in the study of quasinormal modes [Hatsuda:2021gtn, Noda:2022zgk], the Wronskian approach offers both analytical and numerical efficiency [olver1997asymptotics, ince2012ordinary, slavjanov2000special]. Both of these methods should be linked, as they share the same underlying mathematical structures.
@@ 2.1 Generalized Heun Equation
where $z_{i}$ are the finite singular points, $\{0,1-\gamma_{i}\}$ are the characteristic exponents (monodromy parameters) of the finite singularities $z_{i}$, $\{\alpha,\beta\}$ the characteristic exponents for the singular point at $\infty$ The characteristic exponents $\rho$ are determined as the roots of the indicial equation associated with (2.1), $\rho\,(\rho-1)+b_{i}\,\rho+c_{i}=0\,$, (2.2) $b_{i}=\mathrm{Res}_{z=z_{i}}\sum_{i=1}^{n}\frac{\gamma_{i}}{z-z_{i}}\,,\qquad c_{i}=\mathrm{Res}_{z=z_{i}}(z-z_{i})\sum_{i=1}^{n}\frac{q_{i}}{z-z_{i}}\,$, (2.3) associated to each finite regular singular point. For the singular point at infinity, the coefficients in Eq. (2.2) are replaced as $b_{i}\rightarrow b_{\infty}=-\mathrm{Res}_{z=\infty}\sum_{i=1}^{n}\frac{\gamma_{i}}{z-z_{i}}\,,\qquad c_{i}\rightarrow c_{\infty}=\mathrm{Res}_{z=\infty}z\sum_{i=1}^{n}\frac{q_{i}}{z-z_{i}}\,.$ (2.4) , and $q_{i}$ the accessory parameters. The accessory parameters $q_{i}$, finite regular singular points $z_{i}$ and $z=\infty$ characteristic exponents $\{\alpha,\beta\}$ satisfy the following relations
@@ 2.2 Clines and Differential Equations
In the generalized Heun case, when there are more than four finite singularities, it is generally not possible for all of them to lie on a single circle or line. Hence, the geometric configuration of the singularities becomes more intricate reflecting the increased geometric complexity of the generalized case §§§As noted earlier, generalized Heun equations satisfy the Fuchs relation (2.6). This condition, which arises from the presence of only regular singular points, places them within the so-called class of Fuchsian equations. These equations can be divided into two main categories—reducible and irreducible—based on their covariance under field redefinitions of the dependent variable. In the mathematical literature, such transformations are referred to as s-homotopic transformations. For a detailed formal discussion, see Slavyanov and Lay [slavjanov2000special]... One can still use Möbius transformations to simplify three of the singular points (e.g., mapping them to 0, 11, $\infty$), but the remaining singularities may lie off a single cline. We will now argue that the singularities in black hole perturbation lie on clines, and hence are Moebius invariant. We will later employ this fact to unravel the Myers Perry in 5D and with a precise example in 7D show that the reduce the poles to become a Heun equation.
@@ 4.1 Definition
$R(r)/r^{\hat{\ell}}\rightarrow 1\,,\qquad\text{for}\qquad r\rightarrow\infty\,.$ (4.2)
@@
$R(r)\rightarrow R_{\infty}\,(r^{\hat{\ell}}+{\lambda}_{\hat{\ell}}\,\,r^{-\hat{\ell}-2N})\,$, (4.3)
@@ Acknowledgments
We would like to thank the Mitchell Family Foundation for hosting us during the Cook’s Branch workshop where some of the research was carried out. MJR is partially supported through the NSF grant PHY-2309270. LFT’s work is supported by the USU Presidential Doctoral Research Fellowship, Howard L. Blood Fellowship and Farrell and Ann Edwards Scholarship.

### v2

@@ 1 Introduction
Heun differential equations appear across physics—from quantum mechanics and condensed matter systems to perturbations in curved spacetimes—encoding field dynamics through their singularities and connection structure. In black hole physics, both in four and higher dimensions, Heun equations are particularly relevant to the study of quantum corrections to black hole entropy [Arnaudo:2024bbd, Arnaudo:2025btb], scattering processes and greybody factors [Cvetic:1997uw, Cvetic:2009jn, Jorge:2014kra], quasinormal modes [Berti:2009kk, Fioravanti:2021dce, Hatsuda:2021gtn, Bianchi:2021mft, Noda:2022zgk, Lei:2023mqx, Silva:2025khf, Chen:2025sbz], black hole jets [Camilloni:2023wyn] and tidal Love numbers [Kol:2011vg, Hui:2020xxx, Hui:2021vcv, Hui:2022vbh, Rodriguez:2023xjd, Perry:2023wmm, Charalambous:2023jgq, Perry:2024vwz, Glazer:2024eyi]. Beyond this context, similar techniques have found applications in holography, particularly in the study of renormalization group flows [Caceres:2021fuw, Caceres:2022smh], correlation functions [Gutperle:2025bzr] and deformations of the boundary conformal field theory [Karch:2025hof].
Over the last few years, the Heun connection formula program has achieved extraordinary success in describing black hole perturbations, including scattering processes [Bonelli:2021uvf, Bonelli:2022ten]. This was first implemented to calculate Schwarzschild and Kerr quasinormal modes [Aminov:2020yma]We thank Alba Grassi for clarifying this point to us. The method addresses the connection problem — which involves establishing the matrices that connect Frobenius series solutions around distinct singular points — for the local solutions of the four-point Heun equation, employing the so-called AGT correspondence introduced by Alday, Gaiotto, and Tachikawa [Alday:2009aq]. A central aspect is the use of conformal blocks provided by this duality to construct explicit connection formulas between the local solutions of the Heun equation [Bonelli:2022ten] and for Generalized Heun equations with five poles [Arnaudo:2025kof]. This approach has also been generalized to the confluent limit of the Heun equation [Bonelli:2021uvf, Bonelli:2022ten, Consoli:2022eey, Aminov:2023jve, Jia:2024zes, Gutperle:2025bzr]. An alternative approach for deriving connection formulas is based on expressing them as ratios of Wronskians constructed from independent solutions of the differential equation. This method, grounded in the analytic continuation properties of linear differential equations, allows one to extract connection coefficients without explicitly matching local Frobenius series. As demonstrated by Hatsuda, Kimura, Noda, and Motohashi in the study of quasinormal modes [Hatsuda:2021gtn, Noda:2022zgk], the Wronskian approach offers both analytical and numerical efficiency [olver1997asymptotics, ince2012ordinary, slavjanov2000special]. Both of these methods should be linked, as they share the same underlying mathematical structures.
@@ 2.1 Generalized Heun Equation
where $z_{i}$ are the finite singular points, $\{0,1-\gamma_{i}\}$ are the characteristic exponents (monodromy parameters) of the finite singularities $z_{i}$, $\{\alpha,\beta\}$ the characteristic exponents for the singular point at $\infty$ §§§The characteristic exponents $\rho$ are determined as the roots of the indicial equation associated with (2.1), $\rho\,(\rho-1)+b_{i}\,\rho+c_{i}=0\,$, (2.2) $b_{i}=\mathrm{Res}_{z=z_{i}}\sum_{i=1}^{n}\frac{\gamma_{i}}{z-z_{i}}\,,\qquad c_{i}=\mathrm{Res}_{z=z_{i}}(z-z_{i})\sum_{i=1}^{n}\frac{q_{i}}{z-z_{i}}\,$, (2.3) associated to each finite regular singular point. For the singular point at infinity, the coefficients in Eq. (2.2) are replaced as $b_{i}\rightarrow b_{\infty}=-\mathrm{Res}_{z=\infty}\sum_{i=1}^{n}\frac{\gamma_{i}}{z-z_{i}}\,,\qquad c_{i}\rightarrow c_{\infty}=\mathrm{Res}_{z=\infty}z\sum_{i=1}^{n}\frac{q_{i}}{z-z_{i}}\,.$ (2.4) , and $q_{i}$ the accessory parameters. The accessory parameters $q_{i}$, finite regular singular points $z_{i}$ and $z=\infty$ characteristic exponents $\{\alpha,\beta\}$ satisfy the following relations
@@ 2.2 Clines and Differential Equations
In the generalized Heun case, when there are more than four finite singularities, it is generally not possible for all of them to lie on a single circle or line. Hence, the geometric configuration of the singularities becomes more intricate reflecting the increased geometric complexity of the generalized case As noted earlier, generalized Heun equations satisfy the Fuchs relation (2.6). This condition, which arises from the presence of only regular singular points, places them within the so-called class of Fuchsian equations. These equations can be divided into two main categories—reducible and irreducible—based on their covariance under field redefinitions of the dependent variable. In the mathematical literature, such transformations are referred to as s-homotopic transformations. For a detailed formal discussion, see Slavyanov and Lay [slavjanov2000special]... One can still use Möbius transformations to simplify three of the singular points (e.g., mapping them to 0, 11, $\infty$), but the remaining singularities may lie off a single cline. We will now argue that the singularities in black hole perturbation lie on clines, and hence are Moebius invariant. We will later employ this fact to unravel the Myers Perry in 5D and with a precise example in 7D show that the reduce the poles to become a Heun equation.
@@ 4.1 Definition
$R(r)/r^{\ell}\rightarrow 1\,,\qquad\text{for}\qquad r\rightarrow\infty\,.$ (4.2)
@@
$R(r)\rightarrow R_{\infty}\,(r^{\ell}+{\lambda}_{\hat{\ell},m}\,\,r^{-\ell-2N})\,$, (4.3)
@@ Acknowledgments
We would like to thank Adam R. Solomon and Alba Grassi for useful comments on the present work and related topics. We thank also the Mitchell Family Foundation for hosting us during the Cook’s Branch workshop where some of the research was carried out. MJR is partially supported through the NSF grant PHY-2309270. LFT’s work is supported by the USU Presidential Doctoral Research Fellowship, Howard L. Blood Fellowship and Farrell and Ann Edwards Scholarship.